### v1

@@ 2.2 Gaussian Mixture Approximation
(20) $\begin{split}\dot{m}_{k}&=-C_{k}\int\mathcal{N}_{k}(\theta)\Bigl{(}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\Phi_{R}\Bigr{)}\mathrm{d}\theta,\\ \dot{C}_{k}&=-C_{k}\Bigl{(}\int\mathcal{N}_{k}(\theta)\bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{)}\mathrm{d}\theta\Bigr{)}C_{k},\\ \dot{w}_{k}&=-w_{k}\int\Bigl{(}\mathcal{N}_{k}(\theta)-\rho_{a}^{\rm GM}\Bigr{)}\bigl{(}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta.\end{split}$
Here $\lambda=-\int\rho_{a}^{\rm GM}\bigl{(}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta$. Similar to Eq. 14, the covariance evolution equation can be rewritten as:
(21) $\frac{\mathrm{d}C_{k}^{-1}}{\mathrm{d}t}$ $=\int\mathcal{N}_{k}(\theta)\bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{)}\mathrm{d}\theta.$
And the weight evolution equation can be reformulated as
(22) $\frac{\mathrm{d}\log w_{k}}{\mathrm{d}t}$ $=-\int\Bigl{(}\mathcal{N}_{k}(\theta)-\rho_{a}^{\rm GM}\Bigr{)}\bigl{(}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta$
@@ 3 Derivative Free Gaussian Mixture Variational Inference
In this section, we introduce a stable, derivative free approximation of Eq. 20 for Bayesian inverse problems, termed Derivative Free Gaussian Mixture Variational Inference (DF-GMVI). A key component of practical sampling algorithms is the design of effective quadrature rules. Commonly used methods include the mean-point approximation, unscented transformations [38, 36], cubature transformations [2], the Monte Carlo approach, and stochastic approximation techniqeus [29]. The DF-GMVI method employs two quadrature rules, as defined in Definitions 3.1 and 3.2, based the following considerations:
  1. 1.
    Numerical experiments presented in Appendix A demonstrate that, in this context, the mean-point approximation achieves superior convergence.
  2. 2.
    Building on the mean-point approximation, additional corrections are developed to approximate Hessian expectations, capturing as much curvature information of $\mathcal{F}$ as possible, while maintaining linear complexity in the evaluation of $\mathcal{F}$. These corrections ensure covariance positivity and affine invariance, significantly improving the stability of the overall algorithm, as detailed in Section 4.
@@
$\theta_{0}=m\qquad\theta_{i}=m+\alpha[\sqrt{C}]_{i}\qquad\theta_{N_{\theta}+i}=m-\alpha[\sqrt{C}]_{i}\quad(1\leq i\leq N_{\theta}),$
where $\sqrt{C}$ is the square root matrix of $C$, such that $C=\sqrt{C}\sqrt{C}^{T}$. And $[\sqrt{C}]_{i}$ denotes its $i$-th column. Given $\mathcal{F}:\mathbb{R}^{N_{\theta}}\rightarrow\mathbb{R}^{N_{x}}$, we compute vectors
@@
To reach Definition 3.1, we first define ~$\widetilde{\mathcal{F}}(\widetilde{\theta})=\mathcal{F}(m+\sqrt{C}\widetilde{\theta})$. The expectations of $\Phi_{R}(\theta)$ and its gradient in Eqs. 25 and 26 are then approximated using the mean-point approximation with the finite difference approximation of the gradient $\nabla_{\widetilde{\theta}}\widetilde{\mathcal{F}}(0)^{T}$, as given in Eq. 24. The expectation of the Hessian in Eq. 27 can be interpreted as a correction that incorporates curvature information into the mean-point approximation in the Gaussian Newton formulation, $\sqrt{C}^{-T}(B^{T}B)\sqrt{C}^{-1}$. The derivation of the additional terms $\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A))\sqrt{C}^{-1}$ proceeds as follows: We assume that ~$\widetilde{\mathcal{F}}(\widetilde{\theta})=A\widetilde{\theta}\odot\widetilde{\theta}+B\widetilde{\theta}+c$, where $\theta\odot\theta$ denoting elementwise product. The column vectors of matrices $A$ and $B$ are approximated by finite differences given in Eq. 24. Under this assumption, the expectation of the Hessian has an analytical form:
@@
where $D$ is a diagonal matrix with $D_{jj}=2\sum_{k}a_{j}^{T}a_{k}+4a_{j}^{T}a_{j}$. To ensure covariance positivity, we retain only the positive definite part of this expectation, resulting in Eq. 27.
Moreover, it is worth noticing the quadrature rule approximation Eq. 27 achieves exactness when $\mathcal{F}$ is linear. However, black-box variational inference [52] formulates this term as follows:
$\mathbb{E}_{\mathcal{N}}[\nabla_{\theta}\nabla_{\theta}\Phi_{R}(\theta)]=2\int\nabla_{C}\mathcal{N}(\theta;m,C)\Phi_{R}(\theta)\mathrm{d}\theta=C^{-1}\mathbb{E}_{\mathcal{N}}\Bigl{[}\bigl{(}(\theta-m)(\theta-m)^{T}-C\bigr{)}\Phi_{R}\Bigr{]}C^{-1},$
which is derived using integration by parts. Evaluating this requires computing the expectation of a highly nonlinear matrix function. Even when $\mathcal{F}$ is linear, using either the unscented [38, 36] or cubature [2] transformations does not provide an accurate approximation. And stochastic approximation suffers from high variance, which slows convergence (see Section 5.2).
Then, we develop the quadrature rule for approximating the expectation of $\log\rho^{\rm GM}$ and its derivatives used in Eq. 20.
@@
(28) $\mathbb{E}_{\mathcal{N}_{k}}[\log\rho_{a}^{\rm GM}(\theta)]\approx\log\rho_{a}^{\rm GM}(m_{k})\qquad\mathbb{E}_{\mathcal{N}_{k}}[\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)]\approx\nabla_{\theta}\log\rho_{a}^{\rm GM}(m_{k}).$
@@
(29) $\begin{split}&\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)\Bigr{]}\\ &=-\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\frac{\Bigl{(}\sum_{i}w_{i}v_{i}\mathcal{N}_{i}(\theta)\Bigr{)}\Bigl{(}\sum_{i}w_{i}v_{i}\mathcal{N}_{i}(\theta)\Bigr{)}^{T}}{\rho_{a}^{\rm GM}(\theta)^{2}}+\frac{\sum_{i}w_{i}\Bigl{(}C_{i}^{-1}-v_{i}v_{i}^{T}\Bigr{)}\mathcal{N}_{i}(\theta)}{\rho_{a}^{\rm GM}(\theta)}\Bigr{]}\\ &=\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}-v_{j}\bigr{)}\bigl{(}v_{i}-v_{j}\bigr{)}^{T}\mathcal{N}_{i}(\theta)\mathcal{N}_{j}(\theta)}{\rho_{a}^{\rm GM}(\theta)^{2}}\Bigr{]}-\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\frac{\sum_{i}w_{i}\mathcal{N}_{i}(\theta)C_{i}^{-1}}{\rho_{a}^{\rm GM}(\theta)}\Bigr{]}\\ &\approx\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}^{T}\mathcal{N}_{i}(m_{k})\mathcal{N}_{j}(m_{k})}{\rho_{a}^{\rm GM}(m_{k})^{2}}-C_{k}^{-1}.\end{split}$
@@
The proposed quadrature rule in Definition 3.2 is consistent with Definition 3.1, which is important because the variational inference relies on the balance of these two terms. Specifically, the expectation of $\log\rho_{a}^{\rm GM}(\theta)$ and its gradient are computed using the mean-point approximation. For the Hessian approximation in Eq. 29, it is decomposed into two terms. The first term is approximated using the mean-point approximation. The second term is approximated as $C_{k}^{-1}$, which is a correction to the mean-point approximation. This correction satisfies the following equality:
$\sum_{k}w_{k}\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\frac{\sum_{i}w_{i}\mathcal{N}_{i}(\theta)C_{i}^{-1}}{\rho_{a}^{\rm GM}(\theta)}\Bigr{]}=\int\sum_{i}w_{i}\mathcal{N}_{i}(\theta)C_{i}^{-1}\mathrm{d}\theta=\sum_{k}w_{k}C_{k}^{-1}.$
When $K=1$, this correction becomes exact. This correction ensures covariance positivity, significantly enhancing the stability of the algorithm (see Proposition 4.1).
Finally, we update the covariances, means, and weights sequentially using a forward Euler scheme from time $t$ to time $t+\Delta t$ with time step $\Delta t$. For the mean update, we use the updated covariance $C_{k}(t+\Delta t)$ in a Gauss–Seidel manner on the right-hand side.
(30) $\begin{split}C_{k}^{-1}(t+\Delta t)&=C_{k}^{-1}(t)+\Delta t~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{\}},\\ m_{k}(t+\Delta t)&=m_{k}(t)-\Delta t~{}C_{k}(t+\Delta t)~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\Phi_{R}\bigr{\}},\\ \log w_{k}(t+\Delta t)&=\log w_{k}(t)-\Delta t~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\log\rho_{a}^{\rm GM}(t)+\Phi_{R}\bigr{\}}.\end{split}$
Here, $\mathrm{\mathsf{QR}}$ denotes the quadrature rules as defined in Definitions 3.1 and 3.2, and $\mathcal{N}_{k}$ denotes the $k$-th Gaussian component at time $t$. We then normalize ${w_{k}(t+\Delta t)}_{k=1}^{K}$ and, for efficiency, set a lower bound of $w_{k}$ at a default value of $10^{-8}$ during normalization. The computational cost with respect to evaluating $\mathcal{F}$ is $(2N_{\theta}+1)KN_{t}$. This is because evaluating the quadrature rule in Definition 3.1 requires computing $\mathcal{F}$ at $12N_{\theta}+$ 1 quadrature points for each Gaussian mode, where $N_{t}$ represents the total number of time steps. It is worth mentioning the evaluation of $\mathcal{F}$ at the $(2N_{\theta}+1)K$ quadrature points can be performed in an embarrassingly parallel manner.

### v3

@@ 2.1 Approximate Fisher Information
(6) $\begin{split}\frac{\mathrm{d}m_{k}}{\mathrm{d}t}&=-C_{k}\int\mathcal{N}_{k}(\theta)\Bigl{(}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\Phi_{R}\Bigr{)}\mathrm{d}\theta,\\ \frac{\mathrm{d}C_{k}^{-1}}{\mathrm{d}t}&=\int\mathcal{N}_{k}(\theta)\bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{)}\mathrm{d}\theta,\\ \frac{\mathrm{d}\log w_{k}}{\mathrm{d}t}&=-\int\Bigl{(}\mathcal{N}_{k}(\theta)-\rho_{a}^{\rm GM}\Bigr{)}\bigl{(}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta.\end{split}$
A detailed derivation of Eq. 6, along with its connection to natural gradient descent for Gaussian variational families, is presented in the supplementary material. In the following, we refer to Eq. 6 as the natural gradient flow, although it involves an approximation of the Fisher information matrix.
Notably, Eq. 6 is affine invariant—its behavior remains unchanged under any invertible affine transformation, or equivalently, in coordinate systems related by such transformations (see Proposition 5.3). This invariance is particularly beneficial for improving the efficiency of sampling methods when handling highly anisotropic posterior distributions [31, 23, 14].
2.2 Numerical Challenges for Solving the Gradient Flow
Solving Eq. 6 involves evaluating several Gaussian integrals on the right hand side. In the pioneering work [50], Monte Carlo integration was employed to compute these Gaussian integrals, which require both the gradient and Hessian of $\Phi_{R}$. However, to ensure numerical stability, very small time steps were necessary. A subsequent study [15] proposed an efficient, derivative free approximation. In this approach, Monte Carlo integration was applied only to the Gaussian integrals associated with the logarithm of the Gaussian mixture, $\log\rho_{a}^{\rm GM}$, as these functions do not involve the evaluation of forward models, are relatively inexpensive to compute, and admit closed-form expressions. For the integrals involving $\Phi_{R}$, a Kalman-based approach, specifically the unscented transformation [44, 39], was applied. This strategy preserved covariance positivity, enabling significantly larger time steps and resulting in an efficient, derivative free posterior approximation method known as Gaussian Mixture Kalman Inversion (GMKI).
Although GMKI effectively captures multiple modes in the posterior, it struggles to accurately represent the density in regions where different modes significantly overlap (see [15, Section 6.1, Appendix E]). We identify this limitation as resulting from the inconsistent approximation of the two terms: $\log\rho_{a}^{\rm GM}$ and $\Phi_{R}$, which we refer to as the entropy and cross-entropy terms because of their origin in (4). Since the stationary point of Eq. 6 requires a balance between these terms, consistent treatment is important. This issue may be connected in spirit to the phenomenon pointed out by [73], where inconsistency in temporal discretization between the entropy and cross-entropy terms was argued to introduce asymptotic bias when discretizing Langevin dynamics in time.
In the present work, we focus on consistent spatial approximation of Eq. 6 through quadrature. To do so, in the following section, we first consider some numerical examples to investigate the numerical issues in detail. These investigations motivate our practical guidelines for designing accurate and efficient solution schemes.
3 Numerical Motivations and Practical Guidelines
We discuss practical guidelines for the numerical solution of the natural gradient flow Eq. 6, specifically about the parametric approximation and Gaussian integrals. To this end, we consider illustrative 2D examples, which include a 3-mode Gaussian mixture target density, with $\Phi_{R}(\theta)=-\log\Bigl{(}\sum_{i=1}^{3}w^{*}_{i}\mathcal{N}(\theta;m^{*}_{i},C^{*}_{i})\Bigr{)},$ where $w^{*}_{1}=0.3,\ m^{*}_{1}=\begin{bmatrix}1\\ 2\end{bmatrix},\ C^{*}_{1}=I/4$, $w^{*}_{2}=0.4,\ m^{*}_{2}=\begin{bmatrix}2\\ 1\end{bmatrix},\ C^{*}_{2}=I/4$, and $w^{*}_{3}=0.3,\ m^{*}_{3}=\begin{bmatrix}-1\\ -1\end{bmatrix},\ C^{*}_{3}=I/4$, and a circular-shaped target density given by $\Phi_{R}(\theta)=\frac{(1-\lVert\theta\rVert^{2})^{2}}{2\sigma_{\eta}^{2}}$ with $\sigma_{\eta}=0.3$, as described in [15, Appendix E].
We use Gaussian mixtures with $K=10$ components as the parametric family and solve the natural gradient flow Eq. 6 using the forward Euler scheme with a fixed time step size. For this investigation, we will assume access to both the gradient and Hessian of $\log\rho_{a}$ and $\Phi_{R}$. The Gaussian integrals in Eq. 6 are evaluated using various quadrature rules, including the mean-point approximation:
(7) $\int\mathcal{N}(\theta;m,C)g(\theta)\mathrm{d}\theta\approx\int\mathcal{N}(\theta;m,C)\Bigl{(}g(m)+\nabla g(m)(\theta-m)\Bigr{)}\mathrm{d}\theta=g(m),$
where $g$ may be a scalar-, vector-, or matrix-valued function. This corresponds to the linear approximation used in the extend Kalman filter [41]. Other approaches include the unscented transformation [46, 44], which employs a specialized quadrature rule with $12N_{\theta}+$ 1 sigma points, and the Monte Carlo method [35] with $J=20$ ensembles.
We consider four combinations for the choices of quadratures: (1) all $\Phi_{R}$ and $\log\rho_{a}$ related terms are integrated with mean-point approximation, (2) all terms are integrated with unscented transformation, (3) all terms are integrated with Monte Carlo method, and (4) $\Phi_{R}$ related terms are integrated with mean-point approximation, while $\log\rho_{a}$ related terms are integrated with the Monte Carlo method. We note that the fourth combination involves an inconsistent treatment of the integrals.
Figure 1: Quadrature rule comparison for Gaussian mixture and circular-shaped target densities. Each panel displays the reference density and the estimates from approximating the natural gradient flow Eq. 6 with different quadrature rule combinations (left to right). Red circles represent the means $m_{k}$ of each Gaussian component, with gray crosses marking their initial positions. The final panel shows the total variation distance between the reference density and the estimated densities at each iteration.
The results are presented in Fig. 1. We highlight three observations. First, the fifth panels in both tests demonstrate the drawback of the inconsistent approximation of the integrals with respect to $\log\rho_{a}^{\rm GM}$ and $\Phi_{R}$. This combination fails to accurately approximate the delicate regions where different modes significantly overlap. In contrast, a consistent treatment leads to better performance. In fact, such consistent treatment ensures that replacing $\log\rho_{a}^{\rm GM}$ with the exact solution makes the right hand side of Eq. 6 (after the quadrature is applied) equal to zero exactly.
Second, we note that the maximum time steps for stable simulation in these two examples are $3\times 10^{-2}$ and $6\times 10^{-3}$, respectively. For larger time steps, the covariance matrix update loses positivity. For more challenging problems (e.g. Case D in Section 6.2), the maximum stable time step can drop to $10^{-7}$. The positivity of the covariance thus serves as an indicator of stable simulation.
Finally, we observe that the mean-point approximation (7) is the most efficient, as it requires only one function evaluation to compute the integral and performs surprisingly well compared to other quadrature rules. When this approximation is applied, at the stationary point of the method, we have the following conditions:
(8) $\nabla_{\theta}\log\rho_{a}^{\rm GM}(m_{k})+\nabla_{\theta}\Phi_{R}(m_{k})=0,$
$\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(m_{k})+\nabla_{\theta}\nabla_{\theta}\Phi_{R}(m_{k})=0,$
$\log\rho_{a}^{\rm GM}(m_{k})+\Phi_{R}(m_{k})=\text{const}.$
These conditions are not the same as the exact one (i.e., $\log\rho_{a}^{\rm GM}(\theta)+\Phi_{R}(\theta)=\text{const}$) due to approximation error. However, these conditions at least guarantee that the exact condition is satisfied at these mean points $\{m_{k}\}$. In contrast, when other quadrature rules are used, the stationary conditions become linear combinations of $\rho_{a}^{\rm GM}$ and $\Phi_{R}$ that satisfy Eq. 8 (e.g., the Monte Carlo quadrature rule imposes $\frac{1}{J}\sum_{j}\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta^{(j)}_{k})+\frac{1}{J}\sum_{j}\nabla_{\theta}\Phi_{R}(\theta^{(j)}_{k})=0$, where $\theta^{(j)}_{k}\sim\mathcal{N}(m_{k},C_{k})$). This can lead to larger variance and might introduce cancellation errors, making it less indicative that the exact condition is satisfied at any point. The third and fourth panels in the second row of Fig. 1 show that the unscented and Monte Carlo approximations perform worse than the mean-point approximation.
These observations suggest the practical guidance of developing quadrature rules that ensure consistency and guarantee covariance positivity based on the mean-point approximation (7).
@@ 4 Derivative Free Gaussian Mixture Variational Inference
In this section, we introduce a stable, derivative free approximation of Eq. 6 for Bayesian inverse problems in Eqs. 2 and 3, termed Derivative Free Gaussian Mixture Variational Inference (DF-GMVI). The DF-GMVI method employs two derivative free quadrature rules, defined in Definitions 4.1 and 4.2, motivated by the guidelines outlined in Section 3. Specifically, building on the consistent mean-point approximation Eq. 7, we incorporate additional corrections to approximate Hessian expectations, capturing as much curvature information of $\mathcal{F}$ as possible, while maintaining linear complexity in the evaluation of $\mathcal{F}$. These corrections ensure covariance positivity and maintain affine invariance, which contribute to the superior stability of the overall algorithm, as detailed in Section 5.
@@
$\theta_{0}=m,\qquad\theta_{i}=m+\alpha[\sqrt{C}]_{i},\qquad\theta_{N_{\theta}+i}=m-\alpha[\sqrt{C}]_{i}\quad(1\leq i\leq N_{\theta}),$
where $\sqrt{C}$ is the square root matrix of $C$, such that $C=\sqrt{C}\sqrt{C}^{T}$ and $[\sqrt{C}]_{i}$ denotes its $i$-th column. Given $\mathcal{F}:\mathbb{R}^{N_{\theta}}\rightarrow\mathbb{R}^{N_{x}}$, we compute vectors
@@
To derive Definition 4.1, we first define ~$\widetilde{\mathcal{F}}(\widetilde{\theta})=\mathcal{F}(m+\sqrt{C}\widetilde{\theta})$. The expectations of $\Phi_{R}(\theta)$ in Eq. 11 and its gradient in Eq. 12 are then approximated using the mean-point approximation with the finite difference approximation of the gradient $\nabla_{\widetilde{\theta}}\widetilde{\mathcal{F}}(0)^{T}$, as given in Eq. 10. The expectation of the Hessian in Eq. 13 can be interpreted as a correction that incorporates curvature information into the mean-point approximation in the Gaussian Newton formulation, $\sqrt{C}^{-T}(B^{T}B)\sqrt{C}^{-1}$. The derivation of the additional terms $\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A))\sqrt{C}^{-1}$ proceeds as follows: We assume that ~$\widetilde{\mathcal{F}}(\widetilde{\theta})=A\widetilde{\theta}\odot\widetilde{\theta}+B\widetilde{\theta}+c$, where $\theta\odot\theta$ denotes the Hadamard product. The column vectors of matrices $A$ and $B$ are approximated by finite differences given in Eq. 10. Under this assumption, the expectation of the Hessian has an analytical form:
@@
where $D$ is a diagonal matrix with $D_{jj}=2\sum_{k}a_{j}^{T}a_{k}+4a_{j}^{T}a_{j}$. To ensure covariance positivity (see Proposition 5.1), we retain only the positive definite part of this expectation, resulting in Eq. 13.
a comparison, BBVI [63] employs integration by parts to express the terms as follows:
(14a) $\mathbb{E}_{\mathcal{N}}[\nabla_{\theta}\Phi_{R}(\theta)]=C^{-1}\mathbb{E}_{\mathcal{N}}\bigl{[}(\theta-m)\Phi_{R}\bigr{]},$
(14b) $\mathbb{E}_{\mathcal{N}}[\nabla_{\theta}\nabla_{\theta}\Phi_{R}(\theta)]=C^{-1}\mathbb{E}_{\mathcal{N}}\Bigl{[}\bigl{(}(\theta-m)(\theta-m)^{T}-C\bigr{)}\Phi_{R}\Bigr{]}C^{-1}.$
BBVI applies Monte Carlo approximations of the above two expectations, which often lead to high variance.
On the other hand, applying the unscented transformation [46, 44] for evaluating Eq. 14a also gives a finite difference approximation of $\nabla_{\theta}\Phi_{R}(m)$ as in Definition 4.1. However, Eq. 14b involves high-order matrix-valued functions, and directly applying unscented [46, 44], cubature [3] transformations, or Monte Carlo result in high-variance or non-positive estimates which are less accurate or stable numerically.
Then, we discuss the quadrature rules for the term involving $\log\rho_{a}^{\rm GM}$. Note that the gradient and Hessian of $\log\rho_{a}$ are accessible and inexpensive to compute here. We develop quadrature rules to approximate the expectation of $\log\rho_{a}^{\rm GM}$ and its derivatives used in Eq. 6, following the guidelines discussed in Section 3, with particular emphasis on ensuring the consistency with Definition 4.1.
@@
(15) $\mathbb{E}_{\mathcal{N}_{k}}[\log\rho_{a}^{\rm GM}(\theta)]\approx\log\rho_{a}^{\rm GM}(m_{k}),\qquad\mathbb{E}_{\mathcal{N}_{k}}[\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)]\approx\nabla_{\theta}\log\rho_{a}^{\rm GM}(m_{k}).$
@@
[…]
@@
The expectations of $\log\rho_{a}^{\rm GM}(\theta)$ and its gradient in Eq. 15 are computed using the mean-point approximation (8), which is consistent with the approximation used for $\Phi_{R}$ in Definition 4.1. For the Hessian approximation in Eq. 16, it is decomposed into two terms. The first term is positive definite and approximated using the mean-point approximation.The second term is negative definite and may cause a loss of covariance positivity in the update, if we simply use a mean-point approximation. We consider the weighted average which leads to the following equality
$-\sum_{k}w_{k}\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\frac{\sum_{i}w_{i}\mathcal{N}_{i}(\theta)C_{i}^{-1}}{\rho_{a}^{\rm GM}(\theta)}\Bigr{]}=-\int\sum_{i}w_{i}\mathcal{N}_{i}(\theta)C_{i}^{-1}\mathrm{d}\theta=-\sum_{k}w_{k}C_{k}^{-1}.$
This shows that the weighted average of the second term is the same as the weighted average of $-C_{k}^{-1}$. Based on this fact, we approximate the second term as $-C_{k}^{-1}$, rather than using a direct mean-point approximation. This approximation can be understood as a correction. When $K=1$, this approximation becomes exact. Moreover, we will show later in Proposition 5.1 that this correction ensures covariance positivity, significantly enhancing the stability of the algorithm.
With the above quadrature rules, we update the covariances, means, and weights sequentially using a forward Euler scheme from time $t$ to time $t+\Delta t$ with time step $\Delta t$. For the mean update, we use the updated covariance $C_{k}(t+\Delta t)$ in a Gauss–Seidel manner on the right hand side.
(17a) $C_{k}^{-1}(t+\Delta t)$ $=C_{k}^{-1}(t)+\Delta t~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{\}},$
(17b) $m_{k}(t+\Delta t)$ $=m_{k}(t)-\Delta t~{}C_{k}(t+\Delta t)~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\Phi_{R}\bigr{\}},$
(17c) $\log w_{k}(t+\Delta t)$ $=\log w_{k}(t)-\Delta t~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\log\rho_{a}^{\rm GM}(t)+\Phi_{R}\bigr{\}}.$
Here, $\mathrm{\mathsf{QR}}$ denotes the quadrature rules as defined in Definitions 4.1 and 4.2, and $\mathcal{N}_{k}$ denotes the $k$-th Gaussian component at time $t$. We then normalize ${w_{k}(t+\Delta t)}_{k=1}^{K}$ and, for efficiency, set a lower bound of $w_{k}$ at a default value of $10^{-8}$ during normalization. The lower bound, combined with the logarithmic update rule for weights in Eq. 17c, allows weights to recover exponentially from small values, which is advantageous for alleviating potential mode collapse. The computational cost with respect to evaluating $\mathcal{F}$ is $(2N_{\theta}+1)KN_{t}$. This is because evaluating the quadrature rule in Definition 4.1 requires computing $\mathcal{F}$ at $12N_{\theta}+$ 1 quadrature points for each Gaussian mode, where $N_{t}$ represents the total number of time steps. It is worth mentioning the evaluation of $\mathcal{F}$ at the $(2N_{\theta}+1)K$ quadrature points can be performed in an embarrassingly parallel manner.